# MemANNS: Enhancing Billion-Scale ANNS Efficiency with Practical PIM Hardware

### v1

@@ Abstract.
In numerous production environments, including information retrieval and recommendation systems, Approximate Nearest Neighbor Search (ANNS) plays an indispensable role, particularly when dealing with massive datasets that can contain billions of entries. The necessity for rapid response times in these applications makes the efficiency of ANNS algorithms crucial. However, traditional ANNS approaches encounter substantial challenges at the billion-scale level. CPU-based methods are hindered by the limitations of memory bandwidth, while GPU-based methods struggle with memory capacity and resource utilization efficiency.
This paper introduces MemANNS, an innovative framework that utilizes UPMEM’s Processing-in-Memory (PIM) architecture to address the memory bottlenecks in ANNS algorithms at scale. We concentrate on optimizing a well-known ANNS algorithm, IVFPQ, for PIM hardware through several techniques. First, we introduce an architecture-aware strategy for data placement and query scheduling that ensures an even distribution of workload across PIM chips, thereby maximizing the use of aggregated memory bandwidth. Additionally, we have developed an efficient thread scheduling mechanism that capitalizes on PIM’s multi-threading capabilities and enhances memory management to boost cache efficiency. Moreover, we have recognized that real-world datasets often feature vectors with frequently co-occurring items. To address this, we propose a novel encoding method for IVFPQ that minimizes memory accesses during query processing.
Our comprehensive evaluation using actual PIM hardware and real-world datasets at the billion-scale, show that MemANNS offers a significant 4.3x increase in Queries Per Second (QPS) over CPU-based Faiss, and it matches the performance of GPU-based Faiss implementations. Additionally, MemANNS improves energy efficiency, with a 2.3x enhancement in QPS/Watt compared to GPU solutions. These findings suggest that MemANNS can achieve superior performance within a comparable power envelope, highlighting its suitability for large-scale applications such as serving large models.
@@ 1. Introduction
Approximate Nearest Neighbor Search (ANNS) is widely used in many production systems, including data mining (Tagami, 2017), information retrieval (Zhang et al., 2022; Kung et al., 2024), and recommendation systems (Li et al., 2024; Wang et al., 2023; Chen et al., 2022). In these systems, the data typically reaches into the billions. For example, modern social network platforms serve billions of daily active users and provide real-time recommendations (Zhai et al., 2024). Further, these systems frequently operate under stringent latency constraints (e.g., real-time recommendations and interactive user serving), which underscores the critical importance of efficient ANNS algorithms.
ANNS algorithms are broadly classified into four main categories: hash-based (Andoni et al., 2015; Zheng et al., 2016; Gong et al., 2020), tree-based (Chen et al., 2023; Zheng et al., 2020; Arora et al., 2018), graph-based (Kim et al., 2023; Manohar et al., 2024; Wang et al., 2024) and compression-based (Johnson et al., 2019; Liu et al., 2024; Jiang et al., 2023). While graph-based, tree-based and hash-based methods excel in handling million-scale searches, they encounter scalability challenges for billion-scale datasets due to their substantial memory demands. For example, HNSW (Malkov and Yashunin, 2018), a graph-based method, consumes between 60-450 bytes of memory per vertex, resulting in up to 450GB data size for a graph with one billion vertices. In contrast, compression-based methods enable the encoding of extensive datasets into smaller storage footprints, which greatly enhances scalability and cost-efficiency. This is particularly beneficial for applications such as personalized recommendation and large model serving, where managing vast amounts of data is paramount. Thus, in this paper, we focus mainly on compression-based ANNS algorithms. Specifically, we focus on IVFPQ (Jegou et al., 2010), a popular compression-based algorithm known for its effectiveness in reducing the storage of datasets while maintaining a high level of search accuracy.
@@
Figure 1. Query processing time breakdown of the IVFPQ (Jegou et al., 2010) algorithm, using SIFT dataset (Jégou et al., 2011) at different scales (1M, 100M and 1B). Hardware spec. see Section 5.
IVFPQ answers nearest neighbor queries using four stages, namely cluster filtering, lookup table (LUT) construction, distance calculation and top-k identification (see Section 2.1 for details). We study the performance bottleneck of existing IVFPQ implementations (Facebook AI Research, [n. d.]) (CPU-based and GPU-based) and present their runtime breakdown in Figure 1. We adopt datasets with different sizes, from 1 million to 1 billion. We have two interesting observations. First, when data scales to billions, the performance bottleneck shifts significantly compared to smaller datasets. For example, when data scales from 1 million to 1 billion, the performance bottleneck of the CPU-based implementation shifts from the LUT construction stage, which is mainly compute-intensive, to the distance calculation stage, which is mainly memory-bound. This is because the growing data size results in increased memory accesses to the encoded points, hence longer memory access time. Second, traditional IVFPQ implementations are facing different performance bottlenecks: at the billion scales, the CPU-based implementation is bounded by memory, while the GPU-based implementation is bounded by the top-k identification stage. This stage has limited parallelism, preventing full utilization of GPU resources and causing prolonged execution time.
The above observations reveal the limitations of existing ANNS solutions. First, due to the shifted performance bound at different scales, existing performance optimizations for million-scale ANNS (Jiang et al., 2023; Abdelhadi et al., 2019) may not be suitable to billion-scale. Second, traditional architectures have distinct limitations that result in performance bottlenecks in ANNS algorithms. That is, the CPU-based solutions suffer from low memory bandwidth, leading to memory bottlenecks, while the GPU-based solutions face low resource utilization, resulting in suboptimal performance and cost-efficiency. In this paper, we propose to adopt a new hardware architecture, namely the UPMEM Processing-in-Memory (PIM) hardware (upm, 2023a), to address the limitations of existing solutions.
@@
UPMEM PIM is a standard DIMM device that integrates processing cores directly into memory chips, enabling in-memory computation to reduce data movement and enhance efficiency (Devaux, 2019). It features multi-threaded DPU cores with direct access to memory, significantly improving performance for data-intensive applications. While recent studies have explored the use of PIM/near-memory technology for ANNS, they are either designed for million-scale (Xu et al., 2023) or rely on simulated architectures (Lee et al., 2022; Liang et al., 2024). To the best of our knowledge, this is the first study to leverage practical PIM hardware to enhance billion-scale ANNS performance.
To fully utilize the computation and bandwidth resources provided by the UPMEM hardware, we face the following challenges. First, UPMEM PIM shows a high aggregated memory bandwidth, relying on the fact that the workload is balanced across all chips on the PIM hardware. However, achieving workload balance is difficult due to the diversity in data sizes and access frequencies of real-world datasets. Second, UPMEM’s unique hardware architecture requires a careful redesign of the ANN search process and efficient memory management strategy to leverage the full potential of PIM computing power and memory bandwidth. Third, although PIM has shown potential for accelerating memory-intensive applications (Chen et al., 2024), its hardware resources remain relatively limited compared to traditional CPUs and GPUs. This necessitates the design of pruning methods to reduce the workloads for PIM while preserving query accuracy.
@@
In this paper, we propose MemANNS, an efficient framework for ANNS algorithms using UPMEM PIM hardware. MemANNS incorporates several techniques to address the above mentioned challenges individually. First, we introduce an architecture-aware data placement and query scheduling approach to achieve workload balance and maximize the aggregated memory bandwidth of available PIM chips (Section 4.1). Second, we design an efficient thread scheduling approach to fully utilize the multi-threaded computing power on PIM, together with efficient memory management to maximize cache reuse on PIM to further improve the performance (Section 4.2). Third, on observing the fact that in real-world datasets, there are often some items that appear together in data point vectors (Ye et al., 2023). We propose to redesign the encoding in IVFPQ and pre-store the frequently accessed item combinations at offline to reduce the memory accesses at online phase, hence reducing query answering time (Section 4.3). Lastly, we optimize the top-k identification stage with thread-local top-k heap and early pruning to reduce the selection time (Section 4.4). Note that although most of our discussions surround IVFPQ, the framework can also be applied to other quantization-based ANNS algorithms.
We evaluated the performance of MemANNS using 7 UPMEM PIM modules and real-world billion-scale datasets. Experiments have shown that, MemANNS can improve the Query per Second (QPS) by 4.3x compared to Faiss (Facebook AI Research, [n. d.]) on CPU, and is comparable to GPU-based Faiss implementation. When comparing the cost-effectiveness, MemANNS improves the QPS/Watt by 2.3x compared to GPU-based Faiss. That means, if given the same power budget, MemANNS can achieve much better performance than GPU-based ANNS solution. In addition, MemANNS demonstrates near-linear scalability for large-scale datasets, which makes it a promising solution for practical applications such as large model serving (Liang et al., 2024).
@@ 2. Background and Motivation
2.1. Approximate Nearest Neighbor Search (ANNS)
@@
ANNS algorithms are commonly employed to identify the top-k most similar vectors to a query vector within a given set of vectors. Among the various ANNS algorithms available, our study focuses on the Inverted File with Product Quantization (IVFPQ) method (Jegou et al., 2010), which is one of the most popular algorithms for ANNS due to its efficiency in managing large-scale datasets. Figure 2 illustrates a simple example of the IVFPQ process, which can be divided into two phases, namely the offline phase (top) and online phase (bottom).
Offline Phase: The offline phase of the IVFPQ algorithm prepares a dataset for efficient search by creating a compact representation. It employs two key techniques: Inverted File (IVF) and Product Quantization (PQ). IVF clusters search points into $|C|$ clusters using methods like K-means and computes residuals, which are the differences between each point and its cluster centroid. PQ then compresses these residuals by dividing vectors into $M$ subvectors and encoding them using a codebook $B$. Each encoded point in $E$ is an identifier that maps to a specific codeword in the codebook. Practical implementations often use uint8 encoding of an identifier, achieving a compression rate of $4D/M$. This preprocessing enables faster query searches in the online phase.
Online Phase: When a query vector $q$ arrives, the online phase initiates a four-stage process to find the $k$ nearest vectors to $q$. (a) Cluster Filtering: The distances between $q$ and each centroid vector $c$ stored in IVF are calculated and the $nprobs$ closest centroids are selected for further processing. (b) Look-up Table (LUT) Construction: For each selected centroid, an LUT is created to store pre-computed information and eliminate redundant calculations during query searching. Specifically, each element $LUT[j][i]$ represent the distance between the $j$-th vector of sub-codebook $B_{i}$ and the $i$-th sub-segment of the $q-c$ vector. To get the distance from $q$ to a search point $x$, we can look up and accumulate the values in LUTs as: $L2(q,x)=L2(q-c,r)=\sum_{i=0}^{M-1}LUT[{e_{i}}][i]$, where $r$ and $e$ are the residual and encoded vectors of $x$, respectively. (c) Distance Calculation: Using the LUTs, distances from $q$ to all points in the selected clusters are computed by looking up and aggregating the partial distances stored in the LUTs. (d) Identifying Top-K: Lastly, all computed distances are sorted, and the $k$ vectors with the shortest distances are identified as the approximate nearest neighbors.
@@ 2.2. UPMEM PIM Architecture
Processing-in-Memory (PIM) represents a promising paradigm for addressing the data movement bottleneck encountered in many large-scale data analytics applications including ANNS. Recently, real PIM hardware has started to enter the market, with UPMEM PIM being the first commercially available solution (upm, 2023a; Devaux, 2019). It has been applied in recommendation systems (Chen et al., 2024), graph processing (Cai et al., 2024; Giannoula et al., 2024), machine learning (Gogineni et al., 2024; Rhyner et al., 2024) and database systems (Jibril et al., 2024; Bernhardt et al., 2023) to mitigate the memory wall issue. The potential of UPMEM PIM for ANNS algorithms has never been explored.
Figure 3 shows the structure of a UPMEM module, which is a standard DDR4 DIMM module that integrates 16 PIM chips. Each chip encompasses 8 processing cores, referred to as DRAM Processing Units (DPUs). Each DPU is a multi-threaded 32-bit RISC core with a specific ISA (upm, 2023b). This core supports up to 24 hardware threads running at 350 MHz, featuring a 14-stage pipeline. Each DPU has exclusive access to a 64MB DRAM bank named MRAM, a 24KB instruction memory IRAM, and a 64KB scratchpad memory WRAM. DPUs leverage WRAM for processing tasks, and each WRAM access is finished in a single cycle. DPUs have no direct access to MRAM. MRAM accesses are mediated by a DMA engine, with a latency far greater than that of WRAM accesses (Gómez-Luna et al., 2022).
The host CPU interacts with MRAM for data transfer and retrieval. Data transfers between the host CPU and DPUs can happen concurrently if the buffer sizes transferred to/from all MRAM banks are the same; otherwise, the transfers occur sequentially. Notably, inter-DPU communication relies on the host CPU (Hyun et al., 2024). Existing UPMEM PIM systems can support up to 20 DIMMs, each housing 16 chips, enabling a maximum of 2560 DPUs working simultaneously. Given that each DPU has a private connection to its local MRAM, this provides a significant data access bandwidth of up to 7.2 TB/s (Cai et al., 2024), significantly surpassing the memory bandwidth typically available to CPUs.
@@ 2.3. Motivation and Challenges
Opportunities. As shown in Figure 1, traditional computing architectures encounter various performance bottlenecks as data sizes scale to billions in ANNS algorithms. Specifically, for a dataset with one billion vectors, the distance calculation stage, which involves multiple rounds of LUT lookups to retrieve partial distances, requires a total number of $\frac{10^{9}}{|C|}\times nprobs\times M$ memory accesses. For the example in Figure 1, we set $M$ and $nprobe$ to 32 and $|C|$ to 4096, resulting in 250 million random memory accesses per query. This shows significant challenges related to memory bandwidth and latency when deployed on conventional systems. The high aggregated memory bandwidth of the UPMEM PIM architecture (i.e., up to 7.2 TB/s (Cai et al., 2024)) provides an excellent opportunity to mitigate the memory bottleneck and enhance the overall efficiency of ANNS implementations. Although GPUs also provide high memory bandwidth, they often suffer from limited memory capacity. In contrast, the UPMEM PIM offers a scalable solution for managing larger data sizes, thanks to its standard DIMM form factor.
@@
Challenges. The advantage of UPMEM PIM for ANNS stems from its high aggregated memory bandwidth. However, individual DPUs possess restricted storage capacity, computational power, and memory bandwidth. Fully utilizing these resources to create an efficient PIM-based ANNS solution is not easy, particularly with the imbalanced data access pattern and complicated PIM resource management.
Challenge 1: Imbalanced data access pattern. As illustrated in Figure 3, each PIM DIMM consists of 128 DPUs, each with fast and exclusive access to its local memory (i.e., WRAM, IRAM and MRAM). Data communications between DPUs have to go through the host CPU, which is much slower than intra-DPU data access. Thus, it is crucial to have optimized data distribution across DPUs to 1) avoid inter-DPU/CPU-DPU communications as much as possible, and 2) achieve balanced memory accesses across DPUs to maximize the aggregated memory bandwidth. However, achieving this goal is non-trivial, mainly due to the imbalanced data distribution in cluster popularity and cluster size.
We analyzed the access frequencies of different clusters in IVFPQ using the SPACEV1B dataset and found that certain clusters are accessed much more frequently than others (up to 500x) as shown in Figure 4(a). Further, Figure 4(b) illustrates that some clusters contain significantly higher numbers of data points (up to $10^{6}$x). As a result, the two factors lead to much skewed workload across clusters, making it challenging to achieve workload balance across DPUs.
Challenge 2: Complicated PIM resource management. The UPMEM architecture provides parallel structure between DPUs. Each DPU equips with up to 24 hardware threads and a 14-stage pipeline, with exclusive access to its local memory. How to adapt the IVFPQ method to fully utilize the multi-threads and leverage the 14-stage pipeline to maximally reduce the pipeline bubble and hide the memory read latency is a challenging problem. In addition, the fast WRAM on DPU has a limited capacity of 64KB only. Due to the significant bandwidth difference between WRAM and MRAM, we should maximize the amount of data stored in WRAM to improve performance. However, UPMEM lacks a Memory Management Unit (MMU) to virtualize its physical memory. DPUs use physical addresses when accessing memory (Hyun et al., 2024), which prohibits efficient utilization of the limited cache space.
Challenge 3: Limited power of UPMEM PIM. Although UPMEM PIM is equipped with plentiful DPUs (128 per DIMM) and multiple memory layers, its power remains limited in comparison to CPUs with large cores and GPUs boasting thousands of cores. While PIM’s in-memory processing can significantly reduce data movement and associated latencies, the amount of data that can be efficiently processed by these DPUs are constrained. This limitation becomes more pronounced when apply PIM to billion-scale ANNS. This necessitates the design of pruning methods to reduce the workloads for PIM while preserving query accuracy.
In the following section, we give an overview of our MemANNS framework and briefly introduce the techniques proposed in response to these challenges.
Figure 5. Overview of MemANNS. The key optimizations are highlighted in blue.

### v2

# UpANNS: Enhancing Billion-Scale ANNS Efficiency with Real-World PIM Architecture

@@ Abstract.
Approximate Nearest Neighbor Search (ANNS) is a critical component of modern AI systems, such as recommendation engines and retrieval-augmented large language models (RAG-LLMs). However, scaling ANNS to billion-entry datasets exposes critical inefficiencies: CPU-based solutions are bottlenecked by memory bandwidth limitations, while GPU implementations underutilize hardware resources, leading to suboptimal performance and energy consumption. To address these challenges, we introduce UpANNS, a novel framework leveraging Processing-in-Memory (PIM) architecture to accelerate billion-scale ANNS. UpANNS integrates four key innovations, including 1) architecture-aware data placement to minimize latency through workload balancing, 2) dynamic resource management for optimal PIM utilization, 3) co-occurrence optimized encoding to reduce redundant computations, and 4) an early-pruning strategy for efficient top-k selection. Evaluation on commercial UPMEM hardware demonstrates that UpANNS achieves 4.3x higher QPS than CPU-based Faiss, while matching GPU performance with 2.3x greater energy efficiency. Its near-linear scalability ensures practicality for growing datasets, making it ideal for applications like real-time LLM serving and large-scale retrieval systems.
journalyear: 2025conference: The International Conference for High Performance Computing, Networking, Storage and Analysis; November 16–21, 2025; St Louis, MO, USAbooktitle: The International Conference for High Performance Computing, Networking, Storage and Analysis (SC ’25), November 16–21, 2025, St Louis, MO, USAdoi: 10.1145/3712285.3759777isbn: 979-8-4007-1466-5/2025/11
@@ 1. Introduction
Approximate Nearest Neighbor Search (ANNS) is a foundational technology for modern AI systems, enabling applications like real-time recommendations (Li et al., 2024; Wang et al., 2023; Chen et al., 2022) and retrieval-augmented large language models (RAG-LLMs) (Shi et al., 2023; Khandelwal et al., 2019; Jiang et al., 2023b). As datasets grow to billions of entries, which is common in social networks (Zhai et al., 2024) or enterprise-scale LLM serving, efficient ANNS becomes critical.
ANNS algorithms can be categorized into four classes: hash-based (Andoni et al., 2015; Zheng et al., 2016; Gong et al., 2020), tree-based (Chen et al., 2023; Zheng et al., 2020; Arora et al., 2018), graph-based (Kim et al., 2023; Manohar et al., 2024; Wang et al., 2024) and compression-based (Johnson et al., 2019; Liu et al., 2024; Jiang et al., 2023a). While graph-based, tree-based and hash-based methods excel at million-scale searches, they encounter scalability challenges for billion-scale datasets due to prohibitive memory requirements. For example, HNSW (Malkov and Yashunin, 2018), a state-of-the-art graph-based method, requires 60-450 bytes of memory per vertex. This results in a memory requirement of up to 450GB for a billion-vertex graph, making it impractical for real-world deployments. In contrast, compression-based methods dramatically reduce storage footprints by encoding data into compact representations, enabling scalable and cost-efficient billion-scale searches. Thus, this paper focuses mainly on compression-based ANNS algorithms. Specifically, we focus on IVFPQ (Jegou et al., 2010), one of the most popular compression-based algorithms and widely used for accelerating massive industrial systems (e.g., recommendation systems in ByteDance (BytePlus, 2025) and long-context LLM serving (Zhang et al., 2025)).
Limitation of Existing Architectures. IVFPQ integrates two key techniques, including cluster-based filtering (IVF) and product quantization (PQ). IVF partitions data into clusters to narrow search scope and PQ compresses data into compact codes, greatly reducing the storage requirements while maintaining search accuracy. However, our analysis of IVFPQ’s four-stage pipeline, as shown in Figure 1, reveals critical scalability limitations in existing CPU/GPU implementations. When datasets scale from 1M to 1B entries, we identify that: 1) IVFPQ’s runtime bottleneck shifts with dataset scale. For example, when dataset scales from 1M to 1B, the performance bottleneck of CPU implementation shifts from the LUT construction stage, which is mainly compute-intensive, to the distance calculation stage, which is mainly memory-bound. Thus, existing performance optimizations for million-scale ANNS (Jiang et al., 2023a; Abdelhadi et al., 2019) may not be suitable to billion-scale. 2) At billion-scale, both CPU and GPU implementations face various limitations due to architectural mismatch. CPUs become memory bandwidth-limited, struggling to fetch compressed vectors from DRAM for distance calculation. On the other hand, although GPUs excel at parallel distance calculations, they stall during the low-parallelism top-k stage (64% of runtime), causing resource waste.
@@
Figure 1. Query processing time breakdown of the IVFPQ (Jegou et al., 2010) algorithm, using SIFT dataset (Jégou et al., 2011) at different scales (1M, 100M and 1B). Hardware specifications see Section 5.
@@
Opportunities of PIM. To bridge this gap, we adopt a new hardware architecture, namely the UPMEM Processing-in-Memory (PIM) hardware (upm, 2023a), which embed computation within memory to minimize data movement and enhance efficiency (Devaux, 2019). Unlike existing PIM products (Asgari et al., 2021; Lee et al., 2022b; Yuan and Jin, 2025; Tian et al., 2024; Jang et al., 2023), UPMEM PIM is a standard DIMM device that features multi-threaded DPU cores with direct access to memory, significantly improving performance for data-intensive applications. While recent studies have explored the use of PIM/near-memory technology for ANNS, they are either designed for million-scale (Xu et al., 2023) or rely on simulated architectures (Lee et al., 2022a; Liang et al., 2024). To the best of our knowledge, this is the first study to leverage practical PIM hardware to enhance billion-scale ANNS performance.
Challenges. While seem promising, leveraging UPMEM PIM for billion-scale ANNS introduces three challenges: 1 UPMEM’s high aggregated bandwidth depends on evenly distributed workloads across its PIM chips. However, real-world datasets exhibit severe imbalances, skewing compute and memory demands across PIM chips. 2 UPMEM’s unique hardware architecture, e.g., distributed DPU cores, tiered memory, and lack of inter-DPU communication, requires rethinking IVFPQ’s pipeline to maximize efficiency. 3 Despite reducing data movement, UPMEM’s DPUs (350 MHz RISC cores) lack the raw compute power of CPUs/GPUs. This necessitates the design of pruning methods to bypass non-critical computations (e.g., low-impact distance calculations) to maintain low latency while preserving ANN query accuracy.
@@
Innovations. We propose UpANNS, the first framework to enable efficient billion-scale ANNS on commercial PIM hardware. UpANNS integrates four key innovations to address the aforementioned challenges: 1 We introduce architecture-aware data placement and query scheduling to dynamically balance workloads across PIM chips to maximize memory bandwidth utilization (Section 4.1). 2 We design an efficient thread scheduling approach to fully utilize the multi-threaded computing power on PIM, together with efficient memory management to maximize cache reuse on PIM to further improve the performance (Section 4.2). 3 On observing that real-world datasets often contain correlated items in data point vectors, we redesign the encoding in IVFPQ and pre-store frequently accessed item combinations to minimize redundant memory accesses (Section 4.3). 4 We optimize the low-parallelism top-k stage with thread-local top-k heaps and early termination pruning to accelerate the stage on PIM’s limited cores (Section 4.4).
We evaluated the performance of UpANNS using seven UPMEM PIM modules and three real-world billion-scale datasets. Results demonstrates three key advantages over conventional architectures: 1) Performance: UpANNS improves the Query per Second (QPS) by 4.3x compared to CPU-based Faiss (Facebook AI Research, [n. d.]), and is comparable to GPU-based Faiss. 2) Energy efficiency: UpANNS provides 2.3x higher QPS/Watt than GPUs, enabling superior throughput under identical power budgets. 3) Scalability: UpANNS demonstrates near-linear scaling with dataset size, ensuring practical viability for applications like large model serving (Liang et al., 2024).
@@ 2. Background and Motivation
2.1. Approximate Nearest Neighbor Search
@@
Approximate Nearest Neighbor Search (ANNS) algorithms are commonly employed to identify the top-k vectors most similar to a query vector within large datasets. Among the various ANNS algorithms, we focus on the Inverted File with Product Quantization (IVFPQ) method (Jegou et al., 2010), which is one of the most popular ANNS algorithms for its efficiency in billion-scale searches. As shown in Figure 2, IVFPQ operates in two phases, namely the offline preprocessing phase (top) and the online querying phase (bottom).
Offline phase preprocesses dataset to enable rapid searches. First, the Inverted File (IVF) technique partitions data points into $|C|$ clusters (e.g., via K-means), and each point is represented as a residual (i.e., the difference between the point and its centroid). Next, Product Quantization (PQ) compresses these residuals by splitting each vector into $M$ subvectors and encoding them into compact codes using a codebook. For example, a 128-dimensional vector can be reduced from 512 bytes to 64 bytes using uint8 encoding with $M=16$, achieving an 8x compression rate. This compressed representation minimizes storage and accelerates distance computations during queries.
Online phase processes incoming queries in four stages. (a) Cluster filtering computes distances between the query and all cluster centroids, retaining only the nprobenprobe closest clusters to limit the search scope. (b) Lookup table (LUT) construction precomputes partial distances between the query and codebook entries for each subvector, eliminating redundant calculations. (c) During distance calculation, the algorithm aggregates these precomputed LUT values to approximate the distance between the query and every candidate point in the selected clusters. (d) Top-K selection sorts these approximate distances and returns the $k$ nearest neighbors.
@@ 2.2. UPMEM PIM Architecture
Processing-in-Memory (PIM) represents a promising paradigm for addressing the data movement bottleneck encountered in many large-scale data analytics applications including ANNS. Recently, real PIM hardware has started to enter the market, with UPMEM PIM being the first commercially available solution (upm, 2023a; Devaux, 2019). While prior work has demonstrated UPMEM’s utility in recommendation systems (Chen et al., 2024), graph processing (Cai et al., 2024; Giannoula et al., 2024), machine learning (Gogineni et al., 2024; Rhyner et al., 2024) and databases (Jibril et al., 2024; Bernhardt et al., 2023), its potential for ANNS remains unexplored.
A UPMEM module (Figure 3) is a standard DDR4 DIMM housing 16 PIM chips, each containing 8 DRAM Processing Units (DPUs). Each DPU is a 350 MHz RISC core supporting 24 hardware threads and a 14-stage pipeline to hide memory latency (upm, 2023b). The memory hierarchy comprises three tiers, including 1) MRAM (64MB/DPU), which is a slow bulk storage accessed via high-latency DMA transfers; 2) WRAM (64KB/DPU), a fast scratchpad memory with single-cycle access used for active computations and 3) IRAM (24KB/DPU) which is the instruction memory for thread execution.
The host CPU orchestrates data transfers to MRAM, which occur concurrently only if buffer sizes across all DPUs are uniform; otherwise, sequential transfers degrade performance. Critically, DPUs lack direct intercommunication and all coordination must route through the host, introducing overhead for collaborative tasks like top-k aggregation. Despite these constraints, UPMEM’s distributed architecture achieves 20 Tb/s aggregate bandwidth (Cai et al., 2024) by parallelizing access across 2,560 DPUs (20 DIMMs), significantly surpassing the memory bandwidth typically available to CPUs.
@@ 2.3. Motivation and Challenges
Opportunities. Scaling ANNS to billion-entry datasets exposes critical inefficiencies in conventional architectures, as highlighted in Figure 1. Specifically, for a dataset with one billion vectors, the distance calculation stage alone requires 250 million random memory accesses per query (with $M=32$, $|C|=4096$ and $nprobe=32$ as in Figure 1). This overwhelms CPU memory bandwidth (e.g., 85 GB/s) and exceeds GPU memory capacity (e.g., 80GB), creating a prohibitive memory wall. UPMEM PIM directly addresses this bottleneck through its 7.2 TB/s aggregated memory bandwidth (Cai et al., 2024), enabled by parallel access across thousands of DPUs. Unlike GPUs, UPMEM’s standard DIMM form factor supports scalable memory expansion, allowing it to manage datasets far beyond GPU memory limits. By executing computations in-memory, UPMEM eliminates costly data transfers between CPU and memory, making it uniquely suited for memory-intensive ANNS workloads like IVFPQ.
@@
Figure 5. Overview of UpANNS. The grey blocks in each vector represent high frequency co-occurrences in the encoded vectors.
Challenges. While UPMEM PIM offers great potential, realizing its benefits for billion-scale ANNS requires overcoming three key challenges tied to its architecture and workload characteristics:
Challenge 1: Imbalanced data access pattern. IVFPQ’s cluster-centric design interacts poorly with UPMEM’s distributed DPUs. Real-world datasets exhibit extreme skew: popular clusters receive 500x more queries than others (Figure 4(a)), and large clusters contain $10^{6}$x more data points than others (Figure 4(b)). This imbalance creates “hot” DPUs overwhelmed by frequent and large cluster accesses while others idle, wasting UPMEM’s parallel bandwidth. Compounding this, DPUs lack direct communication. Data communications between DPUs have to go through slow host CPU, negating PIM’s latency benefits.
Challenge 2: Complicated PIM resource management. UPMEM provides parallel structure between DPUs. Each DPU equips with 24 hardware threads and 14-stage pipeline, with exclusive access to its local memory. How to adapt IVFPQ to fully utilize the multi-threads and leverage the 14-stage pipeline to maximally reduce the pipeline bubble and hide the memory read latency is a challenging problem. In addition, the fast WRAM on DPU has a limited capacity of 64KB only. Due to the significant bandwidth difference between WRAM and MRAM, we should maximize the amount of data stored in WRAM to improve performance. However, UPMEM lacks a Memory Management Unit (MMU) to virtualize its physical memory. DPUs use physical addresses when accessing memory (Hyun et al., 2024), which prohibits efficient utilization of the limited cache space.
Challenge 3: Limited power of UPMEM PIM. While UPMEM reduces data movement, its DPUs lack the computational density of CPUs/GPUs. Processing millions of memory accesses per query on 350 MHz cores risks prohibitive latency. This necessitates accuracy-aware pruning strategies to skip non-critical computation without degrading query accuracy.